OpenStudy (anonymous):

How to find the derivative of f(t)=t^2-6t-5

OpenStudy (zzr0ck3r):

have you been taught the power rule? \[y = x^n\implies\frac{dy}{dx}=nx^{n-1}\]?

OpenStudy (zzr0ck3r):

n can be rational...

OpenStudy (anonymous):

Yes but for this we are not supposed to use it

OpenStudy (anonymous):

We are supposed to use the normal deriva/ formauls

OpenStudy (zzr0ck3r):

then you need to use \[\lim_{h\rightarrow 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

So do I substitute the equation in for x?

OpenStudy (zzr0ck3r):

im not going to write the limit over and over but \[\frac{f(t+h)-f(t)}{h}=\frac{(t+h)^2-6(t+h)-5}{h}\]

OpenStudy (zzr0ck3r):

can you simplify the numerator of the one on the left?

OpenStudy (zzr0ck3r):

err right*

OpenStudy (anonymous):

Yea okay thank you so much!!

OpenStudy (zzr0ck3r):

let me know what you get at the end so I know its right

OpenStudy (anonymous):

is it 2t+h-5

OpenStudy (zzr0ck3r):

should be 2t-6

OpenStudy (zzr0ck3r):

after you run the limit

OpenStudy (anonymous):

How do you find the limit to sub/ into it

OpenStudy (zzr0ck3r):

we are letting h go to 0 so just make h 0

OpenStudy (zzr0ck3r):

but thats not the problem

OpenStudy (anonymous):

I can't find where I made a mistake

OpenStudy (zzr0ck3r):

sorry I think I messed up the code showing you the first time....

OpenStudy (anonymous):

nvm I found it :( and it was a big wooping stupid mistake, thank you thought!

OpenStudy (zzr0ck3r):

\[\frac{f(x+h)-f(x)}{h}=\frac{(x+h)^2-6(x+h)-5 - (x^2-6x-5)}{h}\\=\frac{x^2+2xh+h^2-6x-6h-5-x^2+6x+5}{h}=\frac{2xh+h^2-6h}{h}=\\2x+h-6\]we let h go to 0 \[=2x+6\]

OpenStudy (zzr0ck3r):

the first think I showed you was wrong...I left off the -f(x) ...woops

OpenStudy (zzr0ck3r):

the first thing I showed you was missing the red\[\frac{f(x+h)-f(x)}{h}=\frac{(x+h)^2-6(x+h)-5 \color{red}{- (x^2-6x-5)}}{h}\]

OpenStudy (anonymous):

Okay cool thank you so much I finally get how to do d/dx